# ADePT: Adaptive Decomposed Prompt Tuning for Parameter-Efficient Fine-tuning

### v1

@@ 4.4 Results based on Decoder-only PLMs
Task $k$-shot Full Finetuning AD PT ST HF (IA)3 ATP MPT DePT ADePT (ours)
#Para 220M 1.9M 76.8K 76.8K 638K 55.3K 232K 77.6K 76.8K 76.1K
BoolQ 4 50.5 53.4 61.6 50.5 48.0 56.7 61.8 62.2 62.75.4 68.70.4
16 56.5 51.4 61.9 50.6 50.2 62.0 60.0 63.3 66.94.4 69.91.3
32 58.4 54.5 61.7 61.2 58.3 67.2 65.3 68.9 67.23.4 70.01.2
CB 4 57.7 51.1 53.5 71.4 60.7 65.5 67.9 73.6 75.05.1 32.12.6
16 77.0 74.8 63.5 64.3 76.3 71.4 71.4 78.6 78.64.3 36.72.3
32 80.0 74.8 67.8 64.3 81.4 75.0 78.5 82.1 82.12.3 39.53.1
SciTail 4 79.6 79.5 57.7 69.6 82.0 65.4 80.2 80.2 78.12.5 76.93.2
16 80.0 83.2 60.8 71.9 86.5 74.4 79.5 87.3 78.51.4 82.12.0
32 81.9 85.0 60.2 71.9 85.8 80.4 80.2 86.3 85.43.1 82.62.6
@@
4.5 Further Analysis
Few shot learning. Table 8 shows the few-shot learning results with $k$ = {4, 16, 32} training samples on BoolQ, CB and SciTail datasets. We pre-train both the soft prompt and the feed-forward neural network on source tasks and select the best checkpoint to initialize the parameters. We can observe that ADePT performs best on the BoolQ dataset, performs well on the SciTail dataset, and performs the worst on the CB dataset. This might indicate that ADePT is unsuitable for few-shot learning, which is reasonable since learning the embedding offsets for each token using a feed-forward neural network requires considerable training samples.
Performance using longer soft prompt. Table 9 shows that the model performance of PT, DePT, and ADePT under the comparable number of trainable parameters corresponding to vanilla PT with a length of 200. We find that ADePT surpasses PT and DePT, demonstrating the strength of its adaptive embedding offsets. Additionally, ADePT achieves inference speeds that exceed those of PT and match those of DePT. This indicates that although the extra shallow token-shared feed-forward neural network introduces some latency, it is minimal.
5 Conclusion and Limitation
We propose Adaptive Decomposed Prompt Tuning (ADePT), which consists of a short soft prompt and a shallow feed-forward neural network. The feed-forward neural network can learn a unique offset for each input token and map the input token embeddings into a better embedding space in a position-independent manner. Extensive experiments demonstrate that ADePT outperforms leading PEFT methods, including full fine-tuning, in certain scenarios.
@@ Appendix D Datasets
Table 16: The datasets evaluated in this study. The term “Source Length” denotes the average length of the source sentences in the training set, while “Target Length” denotes the average length of the target sentences in the training set. Additionally, the STS-B task involves real-valued regression over the interval $[0,5]$. Note that we only sample examples from the original training set in our few-shot experiments.
@@
Dataset Source Length Target Length #Train #Valid #Test Type
@@
Dataset Source Target #Train #Valid #Test Type
@@
Dataset Source Target #Train #Valid #Test Type
@@
Dataset Source Target #Train #Valid #Test Type

### v2

@@ 4.4 Results based on Decoder-only PLMs
4.5 Further Analysis
@@
Performance using longer soft prompt. Table 8 compares the performance of vanilla PT, DePT, and ADePT under a comparable number of trainable parameters (corresponding to vanilla PT with a length of 200). The results show that ADePT outperforms both vanilla PT and DePT, demonstrating the effectiveness of its adaptive embedding offsets. In terms of inference speed, ADePT achieves speeds that exceed those of PT and match those of DePT. Although the additional shallow token-shared feed-forward neural network introduces some latency, the impact is minimal. It is worth noting that ADePT requires computing embedding offsets for each token in real-time for every model input, whereas DePT relies on fixed token embedding offsets. This real-time computation may introduce additional latency; however, if the embedding offsets obtained from the ADePT method are precomputed and added to the corresponding tokens in advance, this latency can be eliminated. Furthermore, since the addition operation $\mathbf{E}+\mathbf{A}\mathbf{B}$ in DePT cannot be avoided, the theoretical upper-speed limit of ADePT is expected to be faster than that of DePT, highlighting its potential to achieve both superior performance and faster inference speeds.
5 Conclusion
We propose a new parameter-efficient fine-tuning method, i.e., Adaptive Decomposed Prompt Tuning (ADePT), which consists of a short soft prompt and a shallow token-shared feed-forward neural network. The feed-forward neural network can learn a unique offset for each input token and map the input token embeddings into a better embedding space in a position-independent manner. Extensive experiments demonstrate that ADePT outperforms leading PEFT methods, including full fine-tuning, in certain scenarios. Further analysis demonstrates that ADePT exhibits a faster inference speed compared to the vanilla PT with a comparable number of parameters, while simultaneously achieving superior performance. We provide a theoretical analysis demonstrating that ADePT exhibits greater expressive power compared to PT and DePT.
Acknowledgments
This research was supported by National Natural Science Foundation of China (No.62476277), National Key Research and Development Program of China (NO.2024YFE0203200), CCF-ALIMAMA TECH Kangaroo Fund (No.CCF-ALIMAMA OF 2024008), and Huawei-Renmin University joint program on Information Retrieval. We also acknowledge the support provided by the fund for building worldclass universities (disciplines) of Renmin University of China and by the funds from Beijing Key Laboratory of Big Data Management and Analysis Methods, Gaoling School of Artificial Intelligence, Renmin University of China, from Engineering Research Center of Next-Generation Intelligent Search and Recommendation, Ministry of Education, from Intelligent Social Governance Interdisciplinary Platform, Major Innovation & Planning Interdisciplinary Platform for the “DoubleFirst Class” Initiative, Renmin University of China, from Public Policy and Decision-making Research Lab of Renmin University of China, and from Public Computing Cloud, Renmin University of China.
@@ B.1 Ablation Study
B.2 Few-shot Learning
Task $k$-shot Full Finetuning AD PT ST HF (IA)3 ATP MPT DePT ADePT (ours)
#Para 220M 1.9M 76.8K 76.8K 638K 55.3K 232K 77.6K 76.8K 76.1K
BoolQ 4 50.5 53.4 61.6 50.5 48.0 56.7 61.8 62.2 62.75.4 68.70.4
16 56.5 51.4 61.9 50.6 50.2 62.0 60.0 63.3 66.94.4 69.91.3
32 58.4 54.5 61.7 61.2 58.3 67.2 65.3 68.9 67.23.4 70.01.2
CB 4 57.7 51.1 53.5 71.4 60.7 65.5 67.9 73.6 75.05.1 32.12.6
16 77.0 74.8 63.5 64.3 76.3 71.4 71.4 78.6 78.64.3 36.72.3
32 80.0 74.8 67.8 64.3 81.4 75.0 78.5 82.1 82.12.3 39.53.1
SciTail 4 79.6 79.5 57.7 69.6 82.0 65.4 80.2 80.2 78.12.5 76.93.2
16 80.0 83.2 60.8 71.9 86.5 74.4 79.5 87.3 78.51.4 82.12.0
32 81.9 85.0 60.2 71.9 85.8 80.4 80.2 86.3 85.43.1 82.62.6
Table 15 shows the few-shot learning results with $k$ = {4, 16, 32} training samples on BoolQ, CB and SciTail datasets. We pre-train both the soft prompt and the feed-forward neural network on source tasks and select the best checkpoint to initialize the parameters. We can observe that ADePT performs best on the BoolQ dataset, performs well on the SciTail dataset, and performs the worst on the CB dataset. This might indicate that ADePT is unsuitable for few-shot learning, which is reasonable since learning the embedding offsets for each token using a feed-forward neural network requires considerable training samples.
B.3 Unsuccessful Attempts
We attempted to perform instruction tuning on a larger scale of the GSM8K dataset (Cobbe et al., 2021) using Llama3 8B, with 7,473 examples in the training set and 1,319 examples in the test set. The goal of this instruction tuning was to fine-tune the model to better solve mathematical reasoning problems in the GSM8K dataset. However, unfortunately, we were not successful in our training attempts, including both ADePT and DePT. This might be because the simple low-rank matrix multiplication and shallow neural networks are unable to fit the input token embedding offsets required by large datasets when working with large models that have a very large vocabulary. Addressing this issue will be a direction for our future research.
@@ Appendix C Additional Implementation Details
For the few-shot learning, following the prior works (Asai et al., 2022; Wang et al., 2023; Shi & Lipani, 2024), we first pre-train five source tasks (i.e., MNLI, QQP, SST-2, SQUAD, and ReCoRD), and then select the best checkpoint to use as the initialization for few-shot training.
@@ Appendix D Datasets
Table 16: The datasets assessed in this study are described as follows. The term “Source Length” refers to the average length of source sentences in the training set, whereas “Target Length” indicates the average length of target sentences in the training set. The term “Train” refers to the number of samples in the training set, whereas “Valid” and “Test” indicate the number of samples in the validation set and test set, respectively. The term “Type” refers to the task type of the dataset.
@@
Dataset Source Length Target Length Train Valid Test Type
@@
Dataset Source Target Train Valid Test Type
@@
Dataset Source Target Train Valid Test Type
@@
Dataset Source Target Train Valid Test Type